The product of two consecutive positive even integers is 360, find the integers
Answer by checkley75(3666)

MY MISTAKE iI MISREAD THE PROBLEM. WEEK EYES I SUSPECT.
X(X+2)=360
X^2+2X-360=0
(X-18)(X+20)=0
X-18=0
X=18 ANSWER FOR THE LOWER NUMBER
18+2=20 FOR THE LARGER NUMBER
PROOF
18*20=360
360=360
SORRY SORRY SORRY FOR THE MISTAKE. ONCE AGAIN, THIS PROVES I'M HUMAN. :=(